# Towards Optimal Distributed Edge Coloring with Small Palettes

### v1

@@ 3 Under which condition can we extend partial $(2\Delta-2)$-colorings?
Our algorithm starts by computing a $(\alpha,\beta)$-clustering $\mathcal{C}$ with respect to some independent set $\mathcal{R}$ (the parameters $\alpha$ and $\beta$ will be fixed when they become relevant) by letting each vertex join the cluster of one of its closest nodes in $\mathcal{R}$. For each cluster $C(r)\in\mathcal{C}$, the central vertex $r\in\mathcal{R}$ saves the BFS-tree $T(r)$ that naturally arises from this clustering. Let $T=\bigcup_{r\in\mathcal{R}}T(r)$ be the forest obtained by joining all those BFS-trees. The edges in $T$ will be colored last by our algorithm. The remaining subgraph $G-E(T)$ has maximum degree at most $\Delta-1$ (as every vertex is part of exactly one such BFS-tree). Hence, we can invoke any regular $(2\Delta-1)$-edge coloring algorithm to get a $(2\Delta-3)$-edge coloring $\varphi$ of $G-E_{T}$. Therefore, the goal of this section is to establish under which conditions we can extend $\varphi$ to a proper $(2\Delta-2)$-edge coloring of the entire graph $G$. As the first step towards this goal we consider the case where the subgraph induced by the uncolored edges forms a simple star graph. In this case, the only forbidden configuration is if all leaves of the star graph are incident to the same set of $\Delta-1$ colors, which can be seen in Figure 1(b).
@@
We construct the auxiliary bipartite graph $B=(E_{T}\cup[2\Delta-2],F)$ as follows: For every edge $e\in E_{T}$ and every color $c\in[2\Delta-2]$ we add the edge $(e,c)$ to $F$ if and only if $c$ is available for $e$ under $\varphi$. Next we observe that any $E_{T}$-saturating matching in $B$ corresponds to an extension of $\varphi$ to $E_{T}$. By Theorem 3.2 such a matching exists if and only if for any subset $S\subseteq E_{T}$ it holds that $\lvert N_{B}(S)\rvert\geq\lvert S\rvert$. Since any edge in $T$ has at least $\Delta-1$ colors available and $\lvert E_{T}\rvert\leq\Delta$, the only set left to check is $S=E_{T}$. For contradiction sake, assume that $\lvert N_{B}(E_{T})\rvert=\Delta-1$ and let $r\in V_{T}$ be the center of the star. This implies that for every $e=rv\in E_{T}$ all colors in $C:=[2\Delta-2]\setminus N_{B}(E_{T})$ are already occupied. Since $\lvert C\rvert=\Delta-1\geq\lvert N_{E}(v)\cap N_{E}(V_{T})\rvert$, this is only possible if $\varphi(N_{E}(v)\cap N_{E}(V_{T}))=C$ for all $e\in E_{T}$. But $\lvert\varphi(N_{E}(V_{T}))\rvert\geq\Delta$, a contradiction!
As the next step we show that this condition generalizes to trees as well. More precisely, we will show that if $\varphi$ assigns at least $\Delta$ distinct colors to the set of edges which are adjacent to at least one vertex at distance $k$ from the root, then there is always an extension of $\varphi$ that assigns at least $\Delta$ distinct colors to the edges adjacent to any vertex at distance $k-1$ from the root.
@@
$V_{k}:=\{e\in V_{T}:\mathrm{dist}_{T}(v,r)=k\}\qquad\text{and}\qquad E_{k}=\{e\in E_{T}:\mathrm{dist}_{T}(e,r)=k\}.$
@@
Let $\ell\in\mathbb{N}$ such that $|\varphi(N_{E}(V_{\ell}))|\geq\Delta$. First we observe that for every $k>0$, each edge $e\in E_{k}$ has exactly one neighboring edge in $E_{k-1}$. Therefore, we can always find at least one available color for every such edge, as long as the set $E_{k-1}$ remains uncolored. Using this observation we can assign arbitrary available colors to all edges in $\bigcup_{k\geq\ell}E_{k}$ by processing these edge sets in sequence, starting with the set with highest index and going through each of these sets in an arbitrary order. However, after we finish coloring $E_{\ell}$ we need to start being more deliberate with our color choices. For any $k<\ell$ we need to ensure that $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$ holds after we colored all edges of $E_{k}$. This is due to the fact that we need $\lvert\varphi(N_{E}(V_{1}))\rvert\geq\Delta$ in order to apply Lemma 3.1, which guarantees that we can complete the coloring on the final layer of edges $E_{0}$ connected to the root. Now we show that given $\lvert\varphi(N_{E}(V_{k+1}))\rvert\geq\Delta$ we can color the edges in $E_{k}$ in such a way that $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$ for any $1\leq k<\ell$. For each $e\in E_{k}$ let $N^{+}(e):=N_{E}(e)\cap N_{E}(V_{k+1})$. Since $\lvert N^{+}(e)\rvert=\Delta-1$ we can always find two edges $e,e^{\prime}\in E_{k}$ such that $\varphi(N^{+}(e))\neq\varphi(N^{+}(e^{\prime}))$. We first assign colors to all edges in $E_{k}$ that are not adjacent to $e$ or $e^{\prime}$. Since we assume $\Delta$ to be sufficiently large, there always exists a vertex $v\in V_{k}$ that is not incident to either $e$ or $e^{\prime}$. Thus, after coloring all $\Delta-1$ edges in $N_{E}(v)\setminus E_{k-1}$ incident at $v$ we already have $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta-1$. If $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$, we are already done. Otherwise, if $\lvert\varphi(N_{E}(V_{k}))\rvert=\Delta-1$, then the set $\varphi(N_{E}(V_{k})\cup(N^{+}(e)\cap N^{+}(e^{\prime})))$ contains at most $32\Delta-$ 3 colors. Therefore, either $e$ or $e^{\prime}$ must have a color available in $[2\Delta-2]\setminus\varphi(N_{E}(V_{k}))$. We pick that color for $e$ or $e^{\prime}$, which ensures that $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$. The remaining edges in $E_{k}$ can then be assigned arbitrary available colors again. Finally, to color the last layer $E_{0}$ containing the edges incident to the root, we can apply Lemma 3.1, since the number of distinct adjacent colors is at least $\Delta$.
@@
Now in order to fulfil this condition, our algorithm assigns two exclusive edges to each cluster. Each cluster may then decide independently, whether it wants to keep the current color or change it to the (yet unused) color $22\Delta-$ 2 for each assigned edge. Hence, the set of all those edges needs to be independent. Therefore we initially compute a maximal matching on $G-E_{T}$ and modify it in a way such that each cluster receives at least two edges. We model this adjustment of $M$ as an instance of hypergraph sinkless orientation, where each vertex proposes to a suitable matching edge in its $2$-hop neighborhood. The core requirement for this problem is that the minimum degree on the vertex side must exceed the maximum rank on the hyperedge side. In other words, each cluster must send more proposals to distinct edges than the number of proposals that any edge receives. The details of this procedure will be explained in Section 4.3. Therefore we need the clusters to either expand quickly or not need such edges by virtue of being already edge-degree-choosable. For the latter case, we observe as soon as a connected component contains one cycle of even length or a vertex of degree less than $\Delta$, it is already edge-degree choosable.
@@
If a cluster does not contain a cycle of even length, we need to make sure that it can satisfy the colorful condition from Lemma 3.4. For that we need to make sure that the cluster contains a sufficient number of leaves. We apply a similar argument to the one used in the analysis of the $\Delta$-vertex coloring algorithm in [22] to show that a connected component that does not contain an even cycle must be expanding exponentially.
Lemma 3.8 (Cluster expansion lemma).
@@ 4.1 High level overview of our reduction to MIS
2:Output: a $(2\Delta-2)$-edge coloring $\varphi$ of $G$.
3:
4:Phase 1: Partition vertices into clusters
5:Compute an MIS $\mathcal{I}$ on the power graph $G^{8}$triangleright Theorem B.10
6:$\mathcal{C},E_{T}\leftarrow$ Cluster$(G,\mathcal{I})$triangleright Algorithm 2
7:Compute a $(2\Delta-3)$-edge coloring $\varphi$ of $G-E_{T}$triangleright Lemma 4.3
8:Phase 2: Assign two exclusive edges to each cluster
9:Compute a maximal matching $M$ of $G-E_{T}$triangleright Lemma 4.5
10:Set up auxiliary hypergraph $H=(\mathcal{C},M)$triangleright Definition 4.9
11:Assign two matching edges to each cluster via HSO on $H$triangleright Theorem B.8
12:Phase 3: Switch colors in order to complete the coloring
@@
16:Extend $\varphi^{\prime}$ to a $(2\Delta-2)$-edge coloring $\varphi^{\star}$ of $G$triangleright Lemma 3.4
@@
We start by computing a maximal independent set $\mathcal{I}$ on the power graph $G^{8}$. Next we partition the vertex set into disjoint clusters. Each vertex joins the cluster of one of its closest MIS-nodes, breaking ties arbitrarily, which leads to a $(4,8)$-clustering $\mathcal{C}$. In addition, we store the BFS-tree $(V,E_{T})$ which arises naturally from this clustering. Computing the MIS on $G^{8}$ ensures that the individual clusters all contain sufficiently many vertices, which will become important later. Next, we remove the union of all BFS-trees from $G$ to obtain a subgraph $G-E_{T}\subseteq G$ of maximum degree at most $\Delta-1$. Hence, it suffices to invoke any $(2\Delta-1)$-edge coloring on $G-E_{T}$ in order to get a $(2\Delta-3)$-edge coloring $\varphi$ of $G-E_{T}$.
@@ 4.2 Phase 1: Partition vertices into clusters
3:
4:$E_{T}\leftarrow\emptyset$
5:for each $v\in V$ in parallel do
6:     $C(v)\leftarrow\bot$
7:     if $v\in\mathcal{I}$ then
8:         $C(v)\leftarrow v$
9:     end if
10:     while $C(v)=\bot$ do
11:         receive messages from neighbors
12:         if $\exists\,u\in N(v):C(u)\neq\bot$ then
13:              $C(v)\leftarrow C(u)$
14:              $E_{T}\leftarrow E_{T}\cup\{uv\}$triangleright break ties arbitrarily
15:              send message $C(v)$ to neighbors
16:         end if
17:     end while
18:end for
@@
First we show that $N^{4}(r)\subseteq C^{-1}(r)$ for all $r\in\mathcal{I}$. Let $r\in\mathcal{I},v\in N^{4}(r)$ and assume for contradiction sake that $C(v)=r^{\prime}\neq r$. Then, $\mathrm{dist}(v,r^{\prime})\leq\mathrm{dist}(v,r)\leq 4$ and therefore $\mathrm{dist}(r,r^{\prime})\leq 8$, contradicting the fact that $r$ and $r^{\prime}$ are independent in $G^{8}$.
For the second property assume that there exist $r\in\mathcal{I}$ and $v\in C^{-1}(r)$ such that $\mathrm{dist}(v,r)>8$. This implies that $\mathrm{dist}(v,r^{\prime})\geq\mathrm{dist}(v,r)>8$ for all $r^{\prime}\in\mathcal{I}$, because otherwise $v$ would have joined the cluster of $r^{\prime}$ instead of $r$. But then $v$ would be independent of all $r^{\prime}\in\mathcal{I}$ on $G^{8}$, contradicting the maximality of $\mathcal{I}$!
@@
We observe that the edge set $E_{T}$ returned by Algorithm 2 satisfies $E_{T}=\bigcup_{r\in\mathcal{I}}E(T(r))$. For every vertex $v\in V$ there is an $r\in\mathcal{R}$ such that $v\in C^{-1}(r)$. Since $T(r)$ is a spanning tree of $C^{-1}(r)$, we get that $v$ is incident to at least one edge in $E(T(r))$. Thus, the maximum degree of $G-E_{T}$ is at most $\Delta-1$. Using the standard reduction from vertex coloring to MIS we can compute a $(2\Delta-3)$-edge coloring of $G-E_{T}$ in $T_{\mathrm{MIS}}(\Delta^{2}\cdot n,4\Delta)$ rounds.

### v2

@@ 3 Under which condition can we extend partial $(2\Delta-2)$-colorings?
The goal of this section is to establish conditions under which we can extend $\varphi$ to a proper $(2\Delta-2)$-edge coloring of the entire graph $G$. As a first step, we examine the case where the subgraph induced by the uncolored edges is a simple star. In this setting, the only forbidden configuration occurs when all the leaves of the star graph share the same set of $\Delta-1$ incident colors, which can be seen in Figure 2(b).
@@
We construct the auxiliary bipartite graph $B=(E_{T}\cup[2\Delta-2],F)$ as follows: For every edge $e\in E_{T}$ and every color $c\in[2\Delta-2]$ we add the edge $(e,c)$ to $F$ if and only if $c$ is available for $e$ under $\varphi$. Observe that any $E_{T}$-saturating matching in $B$ corresponds to an extension of $\varphi$ to $E_{T}$. By Theorem 3.2 such a matching exists if and only if for any subset $S\subseteq E_{T}$ it holds that $\lvert N_{B}(S)\rvert\geq\lvert S\rvert$. Since any edge in $T$ has at least $\Delta-1$ colors available and $\lvert E_{T}\rvert\leq\Delta$, the only set left to check is $S=E_{T}$. For contradiction sake, assume that $\lvert N_{B}(E_{T})\rvert=\Delta-1$ and let $r\in V_{T}$ be the center of the star. This implies that for every $e=rv\in E_{T}$ all colors in $C:=[2\Delta-2]\setminus N_{B}(E_{T})$ are already occupied. Since $\lvert C\rvert=\Delta-1\geq\lvert N_{E}(v)\cap N_{E}(V_{T})\rvert$, this is only possible if $\varphi(N_{E}(v)\cap N_{E}(V_{T}))=C$ for all $e\in E_{T}$ and therefore $\varphi(N_{E}(V_{T}))=C$. But $\lvert\varphi(N_{E}(V_{T}))\rvert\geq\Delta$, a contradiction!
As the next step we show that this condition generalizes to trees as well. More precisely, we will show that if $\varphi$ assigns at least $\Delta$ distinct colors to the edges adjacent to at least one vertex at distance $k$ from the root, then there is always an extension of $\varphi$ that assigns at least $\Delta$ distinct colors to the edges adjacent to any vertex at distance $k-1$ from the root.
@@
$V_{k}:=\{v\in V_{T}:\mathrm{dist}_{T}(v,r)=k\}\qquad\text{and}\qquad E_{k}=\{e\in E_{T}:\mathrm{dist}_{T}(e,r)=k\}.$
@@
(a) The edges with different colors palettes available are dashed.
Figure 4: Extending the coloring to a tree.
@@
Let $\ell\in\mathbb{N}$ such that $|\varphi(N_{E}(V_{\ell}))|\geq\Delta$. First we observe that for every $k>0$, each edge $e\in E_{k}$ has exactly one neighboring edge in $E_{k-1}$. Therefore, we can always find at least one available color for every such edge, as long as the set $E_{k-1}$ remains uncolored. Using this observation we can assign arbitrary available colors to all edges in $\bigcup_{k\geq\ell}E_{k}$ by processing these edge sets in sequence, starting with the set with highest index and going through each of these sets in an arbitrary order. However, after we finish coloring $E_{\ell}$ we need to start being more deliberate with our color choices. For any $k<\ell$ we need to ensure that $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$ holds after we colored all edges of $E_{k}$. This is due to the fact that we need $\lvert\varphi(N_{E}(V_{1}))\rvert\geq\Delta$ in order to apply Lemma 3.1, which guarantees that we can complete the coloring on the final layer of edges $E_{0}$ connected to the root. Now we show that given $\lvert\varphi(N_{E}(V_{k+1}))\rvert\geq\Delta$ we can color the edges in $E_{k}$ in such a way that $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$ for any $1\leq k<\ell$. For each $e\in E_{k}$ let $N^{+}(e):=N_{E}(e)\cap N_{E}(V_{k+1})$. Since $\lvert N^{+}(e)\rvert=\Delta-1$ we can always find two edges $e,e^{\prime}\in E_{k}$ such that $\varphi(N^{+}(e))\neq\varphi(N^{+}(e^{\prime}))$. We first assign colors to all edges in $E_{k}$ that are not adjacent to $e$ or $e^{\prime}$. Since we assume $\Delta$ to be sufficiently large, there always exists a vertex $v\in V_{k}$ that is not incident to either $e$ or $e^{\prime}$. Thus, after coloring all $\Delta-1$ edges in $N_{E}(v)\setminus E_{k-1}$ incident at $v$ we already have $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta-1$. If $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$, we are already done. Otherwise, if $\lvert\varphi(N_{E}(V_{k}))\rvert=\Delta-1$, then the set $\varphi(N_{E}(V_{k})\cup(N^{+}(e)\cap N^{+}(e^{\prime})))$ contains at most $32\Delta-$ 3 colors. Therefore, either $e$ or $e^{\prime}$ must have a color available in $[2\Delta-2]\setminus\varphi(N_{E}(V_{k}))$. We pick that color for $e$ or $e^{\prime}$, which ensures that $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$. The remaining edges in $E_{k}$ can then be assigned arbitrary available colors again. For an illustration of this procedure we refer to Figure 4. Finally, to color the last layer $E_{0}$ containing the edges incident to the root, we can apply Lemma 3.1, since the number of distinct adjacent colors is at least $\Delta$.
@@
Now in order to fulfill this condition, our algorithm assigns two exclusive edges to each cluster. Each cluster may then decide independently, whether it wants to keep the current color or change it to the (yet unused) color $22\Delta-$ 2 for each assigned edge. Hence, the set of all those edges needs to be independent. Therefore we initially compute a maximal matching on $G-E_{T}$ and modify it in a way such that each cluster receives at least two edges. We model this adjustment of $M$ as an instance of hypergraph sinkless orientation, where each vertex proposes to a suitable matching edge in its $2$-hop neighborhood. The core requirement for this problem is that the minimum degree on the vertex side must exceed the maximum rank on the hyperedge side. In other words, each cluster must send more proposals to distinct edges than the number of proposals that any edge receives. The details of this procedure will be explained in Section 4.3. Therefore we need that the cluster expand quickly. This is only the case if the are not a lot of cycles present. Fortunately, this case is easy, since as soon as a connected component contains one cycle of even length or a vertex of degree less than $\Delta$, we can always extend any partial $(2\Delta-2)$ edge coloring to the edges of this connected component.
@@
On the other hand, if a cluster does not contain a cycle of even length, we need to ensure that it can satisfy the colorful condition from Lemma 3.4. To this end, it is crucial that the cluster contains a sufficient number of leaves. We apply a similar argument to the one used in the analysis of the $\Delta$-vertex coloring algorithm in [40] to show that a connected component that does not contain an even cycle must be expanding exponentially.
Lemma 3.8.
@@ 4.1 High level overview of our reduction to MIS
Recall that HSO (Theorem A.8) aims to orient the hyperedges of a hypergraph such that each vertex has at least one outgoing hyperedge. In Phase 2 of Algorithm 1, it is used to rearrange the existing matching, ensuring that each cluster is assigned two exclusive edges.
@@
2:Output: $(2\Delta-2)$-edge coloring $\varphi$ of $G$.
3:Phase 1: Partition vertices into clusters (Section 4.2)
4:Compute an MIS $\mathcal{I}\subseteq V$ on the power graph $G^{8}$triangleright Theorem A.10
5:$\mathcal{C},E_{T}\leftarrow$ Cluster$(G,\mathcal{I})$triangleright Algorithm 2
6:Compute a $(2\Delta-3)$-edge coloring $\varphi$ of $G-E_{T}$triangleright Lemma 4.3
7:Phase 2: Assign two exclusive edges to each cluster (Section 4.3)
8:Compute a maximal matching $M$ of $G-E_{T}$triangleright Lemma 4.5
9:Set up auxiliary hypergraph $H=(\mathcal{C},M)$triangleright Definition 4.9
10:Assign two matching edges to each cluster via HSO on $H$triangleright Theorem A.8
11:Move the assigned edges into the neighborhood of the respective clusters $\triangleright$ Lemma 4.14
12:Phase 3: Switch colors in order to complete the coloring (Section 4.4)
@@
16:Extend $\varphi$ to a $(2\Delta-2)$-edge coloring $\varphi^{\star}$ of $G$triangleright Lemma 3.4
@@
We start by computing a maximal independent set $\mathcal{I}$ on the power graph $G^{8}$. Next we partition the vertex set into disjoint clusters. Each vertex joins the cluster of one of its closest MIS-nodes, breaking ties arbitrarily, but consistently, which leads to a $(4,8)$-clustering $\mathcal{C}$. In addition, we store the BFS-tree $(V,E_{T})$ which arises naturally from this clustering. Computing the MIS on $G^{8}$ ensures that the individual clusters all contain sufficiently many vertices, which will become important later. Next, we remove the union of all BFS-trees from $G$ to obtain a subgraph $G-E_{T}\subseteq G$ of maximum degree at most $\Delta-1$. Hence, it suffices to invoke a greedy $(2\Delta-1)$-edge coloring on $G-E_{T}$ in order to get a $(2\Delta-3)$-edge coloring $\varphi$ of $G-E_{T}$.
@@ 4.2 Phase 1: Partition vertices into clusters
3:$E_{T}\leftarrow\emptyset$
4:for each $v\in V$ in parallel do
5:     $C(v)\leftarrow\bot$
6:     if $v\in\mathcal{I}$ then
7:         $C(v)\leftarrow v$
8:     end if
9:     while $C(v)=\bot$ do
10:         receive messages from neighbors
11:         if $\exists\,u\in N(v):C(u)\neq\bot$ then
12:              $C(v)\leftarrow C(u)$
13:              $E_{T}\leftarrow E_{T}\cup\{uv\}$triangleright break ties arbitrarily
14:              send message $C(v)$ to neighbors
15:         end if
16:     end while
17:end for
@@
First we show that $N^{4}(r)\subseteq C^{-1}(r)$ for all $r\in\mathcal{I}$. Let $r\in\mathcal{I},v\in N^{4}(r)$ and assume for contradiction sake that $C(v)=r^{\prime}\neq r$. Then, $\mathrm{dist}(v,r^{\prime})\leq\mathrm{dist}(v,r)\leq 4$ and therefore $\mathrm{dist}(r,r^{\prime})\leq 8$, contradicting the fact that $r$ and $r^{\prime}$ are independent in $G^{8}$. For the second property assume that there exist $r\in\mathcal{I}$ and $v\in C^{-1}(r)$ such that $\mathrm{dist}(v,r)>8$. This implies that $\mathrm{dist}(v,r^{\prime})\geq\mathrm{dist}(v,r)>8$ for all $r^{\prime}\in\mathcal{I}$, because otherwise $v$ would have joined the cluster of $r^{\prime}$ instead of $r$. But then $v$ would be independent of all $r^{\prime}\in\mathcal{I}$ on $G^{8}$, which contradicts the maximality of $\mathcal{I}$.
@@
We observe that the edge set $E_{T}$ returned by Algorithm 2 satisfies $E_{T}=\bigcup_{r\in\mathcal{I}}E(T(r))$. For every vertex $v\in V$ there is an $r\in\mathcal{R}$ such that $v\in C^{-1}(r)$. Since $T(r)$ is a spanning tree of $C^{-1}(r)$, we get that $v$ is incident to at least one edge in $E(T(r))$. Thus, the maximum degree of $G-E_{T}$ is at most $\Delta-1$. Using the standard reduction from vertex coloring to MIS [56] we can compute a $(2\Delta-3)$-edge coloring of $G-E_{T}$ in $T_{\mathrm{MIS}}(\Delta^{2}\cdot n,4\Delta)$ rounds.